OpenStudy (anonymous):

the indefinite integral of (x)e^(x^2)

OpenStudy (anonymous):

\[\int\limits_{}^{}xe ^{x ^{2}}\]\[u=x^2\]\[du=2x\]\[\frac{ du }{ 2 }=xdx\]Now that we have made a substitution, we can now evaluate our new integral with respect to u.\[\frac{ 1 }{ 2 }\int\limits_{}^{}e^udu\]Can you go from there? :)

OpenStudy (anonymous):

can you continue from there please:)

OpenStudy (anonymous):

Do you know \[\int\limits_{}^{}e^udu\]

OpenStudy (anonymous):

e^u right?

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

awesome thank you so much.

OpenStudy (anonymous):

So now plug back in your U and don't forget the 1/2 at the begining ;) Medal? :)

OpenStudy (anonymous):

yepp!

OpenStudy (anonymous):

so how would i solve the integral of xe^x

OpenStudy (anonymous):

Integration by parts